There are a total of 180 students who are part of badminton, table tennis and tennis team of the college. Of which 70% of students play table tennis, 10% play tennis and 60% play badminton. It is also noticed that all students who play tennis also play table tennis but not badminton. Now a student is selected at random, what is the probability that he/she is the part of table tennis team only?

1
Expert's answer
2022-02-16T15:32:42-0500

There are"180(0.1)=18" students who play tennis.

Let "x=" the number of students who play table tennis and badminton.

Then there are "180(0.7)-18-x=108-x" students who play table tennis only.

There are "180(0.6)-x=108-x" students who play badminton only.


"18+108-x+108-x+x=180"

"x=54"

There are"108-54=54" students who play tennis only.

The probability that he/she is the part of table tennis team only is


"P(tennis \ only)=\dfrac{54}{180}=0.3"